Calculate the percentage by mass to volume 0.2 M NaOH solution.
1
Expert's answer
2012-08-23T08:49:18-0400
0.2 M it's 0.2 mol in 1 L
n=m/Mr Mr(molar mass) of naoh =
40
n=m/40
0.2=m/40
m=0.2 * 40 = 8 g

8 g / (1000 g + 8
g)=0.00793
100% * 0.00793=0.793%